# Efficient electronic cooling by niobium-based superconducting tunnel junctions

### v2

@@ Conclusions
In conclusion, we have demonstrated significant cooling of free electrons with respect to lattice phonons up to 2.4 K bath temperature. At 2.4 K, we observe cooling of electrons down to 1.6 K – equivalent to 34% relative cooling. The temperature range was limited by the electron thermometry, i.e., the cooling effect is expected to work similarly also above 2.4 K. The cooling result was enabled by developing Nb-based superconducting tunnel junction refrigerator technology. With our modern foundry compatible Al–AlOx–Nb tunnel junctions, we obtained a cooling power density of mW/mm2, which is expected to enable substantial cooling in practical chip scale refrigerators. Most importantly, after decades of research on tunnel junction based electronic coolers, here we demonstrated a device that can operate at the base temperature of pulse tube refrigerators, thus paving the way for replacing bulky dilution refrigerators by miniaturized electronic coolers that operate without rare and expensive .
@@ Acknowledgements
The research was funded by the European Union’s Horizon RIA, EIC and ECSEL programmes under grant agreements No. 766853 EFINED, No. 824109 European Microkelvin Platform (EMP), No. 101113086 SoCool, No. 101007322 MatQu, and No. 101113983 Qu-Pilot. We also acknowledge financial support of Research Council of Finland through projects No. 322580 ETHEC, No. 356542 SUPSI, the QTF Centre of Excellence project No. 336817, and Business Finland through Quantum Technologies Industrial (QuTI) No. 128291 and Technology Industries of Finland Centennial Foundation.
@@
Research was conceptualized by J.H, A.R, A.K, J.S.L and M.P. Device design was done by J.H, A.R, R.P.L and K.V. Device fabrication was led by R.P.L with the support from J.H. T.R did device validation through SEM imaging. Device characterization was done in room temperature by J.H and J.Ge, and in low temperature by J.H and A.R. Data was analyzed by J.H, A.R, E.M and A.K. Finite element method simulation was done by E.M and A.R. with the support from J.H. Al–AlOx–Nb junction development was done by R.P.L, M.R, J-P.K, O.P, V.V, C.F, E.T.M, M.K, M.P and J.G, who also supervised the junction development. M.P supervised the refrigerator work. J.H and M.P wrote the manuscript with contributions from all other authors.

### v3

@@ Conclusions
In conclusion, we have demonstrated significant cooling of free electrons with respect to lattice phonons up to 2.4-K bath temperature. At 2.4 K, we observe cooling of electrons down to 1.6 K – equivalent to 34% relative cooling. The temperature range was limited by the electron thermometry, i.e., the cooling effect is expected to work similarly also above 2.4 K. The cooling result was enabled by developing Nb-based superconducting tunnel-junction refrigerator technology. With our modern foundry-compatible Al-AlOx-Nb tunnel junctions, we obtained a cooling power density of approximately mW/mm2, which is expected to enable substantial cooling in practical chip scale refrigerators. Most significantly, after decades of research on tunnel-junction-based electronic coolers, here we demonstrated a device that can operate at the base temperature of pulse tube refrigerators, thus paving the way for replacing bulky dilution refrigerators by miniaturized electronic coolers that operate without rare and expensive .
@@ Acknowledgements
The research was funded by the European Union’s Horizon RIA, EIC and ECSEL programmes under grant agreements No. 766853 EFINED, No. 824109 European Microkelvin Platform (EMP), No. 101113086 SoCool, No. 101007322 MatQu, and No. 101113983 Qu-Pilot. We also acknowledge financial support of Research Council of Finland through projects No. 322580 ETHEC, No. 356542 SUPSI, the QTF Centre of Excellence project No. 336817, and Business Finland through Quantum Technologies Industrial (QuTI) No. 128291 and Technology Industries of Finland Centennial Foundation and Chips JU project Arctic No. 101139908.
@@
Research was conceptualized by J.H, A.R, A.K, J.S.L and M.P. Device design was done by J.H, A.R, R.P.L and K.V. Device fabrication was led by R.P.L with the support from J.H. T.R did device validation through SEM imaging. Device characterization was done in room temperature by J.H and J.Ge, and in low temperature by J.H and A.R. Data was analyzed by J.H, A.R, E.M and A.K. Finite-element-method simulation was done by E.M and A.R. with the support from J.H. Al-AlOx-Nb junction development was done by R.P.L, M.R, J-P.K, O.P, V.V, C.F, E.T.M, M.K, M.P and J.G, who also supervised the junction development. M.P supervised the refrigerator work. J.H and M.P wrote the manuscript with contributions from all other authors.